OpenStudy (anonymous):

What is the slope of the line that passes through the points (2, 3) and (–2, 6)?

OpenStudy (anonymous):

k

OpenStudy (anonymous):

same as the way i showed u b4

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

try it out.

OpenStudy (anonymous):

y2-y1/x2-x1

OpenStudy (anonymous):

use the formula y = mx + b if not graph it and count up from you first point to the side to the other, the slope is rise over run

OpenStudy (anonymous):

I think its - 3/4 right

OpenStudy (anonymous):

tell me how u did it plz

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

no its up 3

OpenStudy (anonymous):

it is

OpenStudy (anonymous):

graph it and count

OpenStudy (anonymous):

(2, 3) and (–2, 6) y2-y1/x2-x1= 6-3/-2-2 =-34

OpenStudy (anonymous):

-3/4

OpenStudy (anonymous):

sorry

OpenStudy (anonymous):

3/-4

OpenStudy (anonymous):

but u can write it as -3/4

OpenStudy (anonymous):



OpenStudy (anonymous):

everything is positive, 3/4

OpenStudy (anonymous):

wait no -4 cause it goes left

OpenStudy (anonymous):

oh it is?

OpenStudy (anonymous):

3/-4

OpenStudy (anonymous):

Do you understand @chocolatepie123

OpenStudy (anonymous):

there is no 3/ -4

OpenStudy (anonymous):

what are the choices

OpenStudy (anonymous):

plz write that with the question

OpenStudy (anonymous):

-3/4 there?

OpenStudy (anonymous):

then try going the other way -3/4

OpenStudy (anonymous):

if its not going up use the other ppoint to go down

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Do you understand now? @chocolatepie123

OpenStudy (anonymous):

yea

OpenStudy (anonymous):

good

OpenStudy (anonymous):

good glad i could help :D well @Haiqa did too

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

ur welcome

OpenStudy (anonymous):

who the medal going to

OpenStudy (anonymous):

doesn't matter just glad you understand